OpenStudy (anonymous):

Solve the equation by completing the square: 2x² - 4x = 6

OpenStudy (callisto):

2x² - 4x = 6 x² -2x=3 (x²-2x+1)-1 =3 (x-1)² =4 x-1 =2 or x-1 =-2 x=3 or x=-1

OpenStudy (anonymous):

\[2(x^2-2x)=6\]\[2(x-\frac{2}{2})=6+(1^2*2)\]\[2(x-1)^2=6+2\]\[(x-1)^2=\frac{8}{2}\]\\[x-1=\sqrt{4}\]\[x=1\pm2\]\[x=3,x=-1\]